OpenStudy (anonymous):

Before calculators were around, why was it advantageous to use logarithms to calculate roots?

OpenStudy (anonymous):

because \(\log(\sqrt{a})=\frac{1}{2}\log(a)\) so if you knew \(\log(a)\) you could divide by 2 and then take the "anti log" of the result to find the answer what is this class, ancient history?

OpenStudy (anonymous):

no its Integrated Pre-Calculus Algebra and trigonometry